OpenStudy (anonymous):

Find the center and radius of each circle. Determine the intercepts of the graph of each circle 3x^2+3y^2-6x+12y=0

OpenStudy (mertsj):

Divide by 3 first.

OpenStudy (mertsj):

x^2-2x+y^2+4y=0 Complete the square twice.

OpenStudy (mertsj):

x^2-2x+1 +y^2+4y+4=5 (x-1)^2+(y+2)^2=5 Center = (1,-2) and radius = sqrt5

OpenStudy (mertsj):

Find the x intercepts by letting y = 0 3x^2-6x=0 3x(x-2)=0 x = 0 and x = 2 Those are the x intercepts.

OpenStudy (mertsj):

Find the y intercepts by letting x = 0 3y^2+12y=0 3y(y+4)=0 y=0, y=-4 Those are the y intercepts.